OpenStudy (anonymous):

f(3a)=2x(3a)^2-3(3a)+4?

OpenStudy (anonymous):

What's your question?

OpenStudy (anonymous):

What is the answer to that equation?

OpenStudy (anonymous):

That depends on what you're supposed to do with it. If you're supposed to expand it and simplify, we can do that. Let's start by expanding the square: (3a)^2 = 9 a^2 Simplify the first term.

OpenStudy (jdoe0001):

hmm and equation EQUATes something it really has no answer per se your "statement" is just saying that a function equals something else so.. not really a question

OpenStudy (anonymous):

Yes, but what's the simplified version? It is a question, since I don't know what the simplest form is. Anything you ask that you don't personally have an answer to can be and usually is a question. :/

OpenStudy (anonymous):

It sounds like you're just supposed to simplify the equation. Is that correct? If so, let's look at the first term: 2(3a)^2 = ? First, we expand the squared part: (3a)^2 = 9 a^2 So, what is 2*9*a^2?